OpenStudy (issy14):

I'm confused with this partial derivative: z=ye^x/y . Zy=e^x/y But when I do Zx I get xe^x/y and the book says otherwise, can somebody please explain me what I'm doing wrong?

OpenStudy (phi):

is this \[ z= y\ e^{\frac{x}{y} } \]?

OpenStudy (issy14):

No, the book apparently used the product rule and it has the following: -x/y e^x/y + e^x/y

OpenStudy (issy14):

I don't know where they are getting the negative from

OpenStudy (phi):

what is the original expression ?

OpenStudy (issy14):

z=ye^x/y

OpenStudy (phi):

is x/y in the exponent i.e. as I posted up above ?

OpenStudy (issy14):

yes

OpenStudy (phi):

what do you get for \( z_x \) ?

OpenStudy (issy14):

e^x/y

OpenStudy (phi):

yes, that is what I get.

OpenStudy (issy14):

for Zx yes but the problem is for Zy

OpenStudy (phi):

for Zy, treat x as a constant for y exp( x/y) we need to use the product rule y d/dy exp(x/y) + exp(x/y) d/dy y the 2nd term is just exp(x/y) the first term is y * exp(x/y) * d/dy (x/y)

OpenStudy (issy14):

I get the second term as just being e^x/y. that's fine, I"m confused as to where they are getting the negative from

OpenStudy (phi):

d/dy (x/y) (with x constant ) is x * d/dy y^(-1)

OpenStudy (issy14):

hold on let me get the stylus so that I can write it out

OpenStudy (phi):

you don't have to write it out. you just have to work on the first term \[ y \frac{d}{dy}(e^\frac{x}{y} ) \] the derivative of e^u is e^u du so \[ y \frac{d}{dy}(e^\frac{x}{y} ) \\ y\ e^\frac{x}{y}\frac{d}{dy}\frac{x}{y} \]

OpenStudy (issy14):